OpenStudy (anonymous):

write a quadratic equation with the solutions -8 and 3

OpenStudy (anonymous):

Reverse solve for x. Currently, it's x= -8, 3. Get it to equal 0. x = -8 x + 8 = 0 x = 3 x - 3 = 0 FOIL this to find the answer. (x + 8)(x - 3) = 0